# Ensemble Kalman Filtering-Aided Variational Inference for Gaussian Process State-Space Models

### 23 Dec 2023 v2

@@ IV-C More Discussions and Insights
We next provide a comprehensive rationale for selecting EnKF as the integral component of EnVI. We also present detailed insights into the EnVI and discuss its connections to existing works.
Indeed, there are various well-established model-based filtering and smoothing techniques that can be potentially integrated into the variational inference framework for GPSSMs. Such as KF, EKF, PF, and EnKF, to name a few [2]. Among these options, EnKF, as a Monte Carlo-based method, exhibits superior capability in handling nonlinear systems when compared to KF and EKF [45]. While analytical solutions can be obtained for nonlinear systems using e.g. EKF methods, they do not align with the principle advocated by John Tukey that an approximate solution to the right problem is worth more than a precise solution to the wrong problem [50]. In comparison to PF, EnKF may have certain limitations in dealing with strongly nonlinear and non-Gaussian systems. However, sampling in PF is generally more computationally demanding, particularly for high-dimensional systems [51], whereas EnKF can be effectively employed even in extremely high-dimensional systems [52]. Moreover, the inherent differentiability of each operation within EnKF allows the EnVI to jointly update the parameters ${{\bm{\theta}}}$ and ${\bm{\zeta}}$ through using modern automatic differentiation tools [44]. In contrast, PF involves resampling from discrete distributions [53], which pose challenges for the utilization of the reparameterization trick [48, 54], resulting in significantly higher computational complexity for gradient computations [55].
It is also vital to emphasize that the proposed EnVI algorithm falls under the NMF category (see definition in Appendix C), given the exploitation of the dependencies between latent states and the transition function, as evident in the variational distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ in Eq. (24). This indicates that EnVI inherits the favorable characteristics of NMF algorithms [29], which have the potential to enhance learning accuracy and address the issue of underestimating state inference uncertainty, commonly encountered in MF algorithms [35].
Last but not least, note that research work has been conducted on learning dynamical systems using EnKF, as evidenced by works such as [55, 56, 57] and the references therein. However, it is crucial to emphasize that the proposed EnVI deviates from existing methods in two fundamental aspects.
@@
First, the inherent differentiability of EnVI, spanning from parameters ($\{{{\bm{\theta}}},{\bm{\zeta}}\}$) to latent states and extending to the objective function (the ELBO), enables principled joint learning of dynamics and inference of latent states using automatic differentiation tools such as PyTorch [44]. Most existing methods employ the expectation-maximization (EM) algorithm [11, 58] to iteratively update the model parameters ${{\bm{\theta}}}$ and latent state trajectory ${\vec{\mathbf{x}}}$. However, it has been reported that these methods disregard the gradient information from the parameters ${{\bm{\theta}}}$ to the state trajectory ${\vec{\mathbf{x}}}$, which can potentially degrade the model’s learning performance [55, 58]. In contrast, our method draws inspiration from the state-of-the-art EnKF-based dynamical system learning method, autodifferentiable EnKF (AD-EnKF) [55]. By regarding the latent state trajectory as a function of both the model parameters ${{\bm{\theta}}}$ and variational parameters ${\bm{\zeta}}$, and by leveraging reparameterization tricks and automatic differentiation, we can jointly optimize these parameters in a principled way, resulting in enhanced performance.
@@
Second, EnVI offers an effective means of mitigating the risk of overfitting. In the state-of-the-art AD-EnKF [55], the transition function in the SSM is modeled using a deterministic parametric model, specifically a neural network, and the optimization objective function is the logarithm of marginal likelihood of the model, i.e.
@@
However, maximizing this objective function solely w.r.t. the model parameters can easily lead to the model overfitting. In contrast, our method leverages the Bayesian nonparametric model and the variational inference framework to derive the ELBO, as presented in Eq. (27), from which the additional regularization terms for the initial state and state transition function can effectively mitigate the issue of model overfitting. We defer further discussions to Section VI.
V OEnVI: Online Learning and Inference
In this section, we further explore online setting where data is processed sequentially, one sample at a time. It is within this context that the simultaneous inference of states and nonlinear dynamics in GPSSMs presents significant challenges [59, 60, 61, 62]. The good news is that our EnVI approach readily lends itself to online learning scenarios. At each time step $t$, we can naturally maximize the corresponding objective function, denoted as $L_{\ell_{t}}$, given by
@@ V OEnVI: Online Learning and Inference
in terms of both the model parameters and variational parameters. The pseudocode implementation of online EnVI, termed OEnVI, is summarized in Algorithm 2.
Remark 3.
Proposition 2.
The log-likelihood $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ can be expressed as the difference between two fundamental components: the data reconstruction error term, represented by $\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]$, and the KL divergence term between the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$ and the prediction distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$. That is, an alternative form of the objective function for OEnVI can be expressed as
$\quad-\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\operatorname{KL}[q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})]\right]$
Proof.
The proof can be found in Appendix B
The insight in Proposition 2 sheds light on the interplay between data reconstruction and the alignment of filtering and prediction distributions in the EnVI and OEnVI algorithms.
Up to this point, it is worth noting that in contrast to the existing inference network-based variational algorithms [28, 29, 30, 31, 39, 43, 62, 60], OEnVI is a simple and straightforward extension of EnVI, which benefits from eliminating the dependence on the additional parametric variational distributions. Previous works have typically employed inference networks that take the entire sequence of observations ${\mathbf{y}}_{1:T}$ as input, necessitating a significant amount of data for offline training. While it is conceivable to limit the input length of the inference network to a shorter sequence $l$, such as ${\mathbf{y}}_{t-l:t}$, this approach still leads to prolonged training times and higher computational requirements for optimizing the inference network parameters [62]. Moreover, storing historical inputs ${\mathbf{y}}_{t-l:t}$ adds to the storage overhead, making it problematic in situations where historical data duplication and storage are not permissible. In sharp contrast, OEnVI successfully overcomes the aforementioned challenges related to the storage and optimization of inference networks. As a result, it facilitates more efficient learning and inference processes, contributing to improved overall performance. Furthermore, OEnVI offers a more comprehensive objective by simultaneously minimizing the KL divergence, accounting for data reconstruction error balance, and applying regularization to the transition function to mitigate model overfitting. This stands in contrast to the existing approaches that sorely minimize the KL divergence between filtering and prediction distributions in the context of online learning [61].
@@ VI-B System Dynamics Learning
Figure 4: Kink function learning performance against the training iterations. EnVI exhibits faster convergence compared to vGPSSM and VCDT.
@@ VI-C Time Series Data Forecasting
This subsection demonstrates the series prediction performance of the proposed EnVI algorithm on five public real-world system identification datasets [^4] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html, which consist of one-dimensional time series of varying lengths between $296$ to $1024$ data points. In addition to the comparison methods discussed in Section VI-B, EnVI is also compared with several other competitors, including two NMF class methods, PRSSM [29], ODGPSSM [37], and two inference network-based methods, DKF [43] and CO-GPSSM [39], as depicted in Table II. For each method’s training, the first half of the sequence in every dataset is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence, and the latent state dimension, $d_{x}$, is consistently set to $4$ for all datasets. Table II reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
@@
VI-D Online Learning and Inference with OEnVI
Figs. 5a5c depict the true states (in blue) and the latent states (in red) inferred by the three methods. The results clearly indicate that EnVI and SVMC swiftly captured the real state and maintained accuracy, while VJF faced challenges, primarily due to its difficulty in learning the inference network. A detailed comparison with SVMC, as illustrated in Figs. 5d5f, underscores EnVI’s superior accuracy in both inference and prediction of latent states. These empirical findings emphasize OEnVI’s efficacy, particularly in its use of EnKF to approximate the variational distribution, demonstrating its advancement over VJF using inference networks, and SVMC using PF methods.
@@ VI-D Online Learning and Inference with OEnVI
VII Conclusion and Future Work
In this paper, we have introduced EnVI, a novel NMF algorithm tailored for GPSSMs, which integrates EnKF into a variational inference framework. Additionally, we have presented an extended online version, OEnVI. Both algorithms eliminate the necessity for inference networks and shape the variational distribution over latent states through model-based EnKF, enhancing model training efficiency and algorithm robustness compared to existing methods. Empirical experiments conducted on diverse real and synthetic datasets, evaluating filtering, prediction, and dynamics learning performance, unequivocally support the effectiveness of the proposed methods. Looking ahead, our focus is on further reducing the computational complexity of the GP transition function and enhancing its modeling capacity. This effort aims to render the GPSSM compatible with more high-dimensional datasets, aligning it with the capabilities of EnKF and enabling the application of EnVI in more practical, high-dimensional dynamic systems.
Appendix A Proof of Proposition 1
@@
The proof is shown as follows,
@@
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$ (36e)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{1:T}|{\vec{\mathbf{u}}})\right].$ (36f)
@@ Appendix B Proof of Proposition 2
which completes the proof. Here Eq. (37c) is obtained straightforwardly by applying Bayes’ theorem. ∎
@@ D-A Evidence Lower Bound (ELBO)
In Bayesian statistics, the model marginal likelihood $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is a fundamental quantity for model selection and comparison [42]. By maximizing the logarithm of $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ w.r.t. the model parameters ${{\bm{\theta}}}$, the goodness of data fitting and the model complexity are automatically balanced, in accordance with Occam’s razor principle [41]. However, $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is obtained by integrating out all the latent variables $\{{\vec{\mathbf{x}}},{\vec{\mathbf{f}}}\}$ in the joint distribution, see Eq. (15), which is analytically intractable. Thus, the posterior distribution of the latent variables, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})=\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{y}}}|{{\bm{\theta}}})},$ cannot be expressed in a closed-form analytical expression, either. This intractability issue has been addressed in variational Bayesian methods by adopting a variational distribution [11], $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, to approximate the intractable $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$. With the newly introduced variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, we have

### 08 Jan 2024 v3

@@ IV-C More Discussions and Insights
This subsection presents more detailed insights into the proposed EnVI and discusses its connections to existing works.
@@
First of all, the importance of the differentiable nature within EnKF, as mentioned in Remark 1, to the EnVI becomes more apparent. The inherent differentiability in EnVI, spanning from parameters ($\{{{\bm{\theta}}},{\bm{\zeta}}\}$) to latent states and extending to the objective function (the ELBO), enables principled joint learning and inference using modern off-the-shelf automatic differentiation tools (e.g., PyTorch [42]). This contrasts with most existing EnKF-based dynamical systems learning methods (see, e.g., [51, 52] and the references therein), which employ the expectation-maximization (EM) algorithm [10] to iteratively update the model parameters ${{\bm{\theta}}}$ and latent state trajectory ${\vec{\mathbf{x}}}$. It has been reported that such EM-based methods disregard the gradient information from the parameters ${{\bm{\theta}}}$ to the state trajectory ${\vec{\mathbf{x}}}$, which can potentially degrade the learning performance [47, 53]. In contrast, by regarding the latent state trajectory as a function of both the model parameters ${{\bm{\theta}}}$ and variational parameters ${\bm{\zeta}}$, our method can jointly optimize these parameters in a principled way, resulting in enhanced performance.
Second, it is worth noting that the proposed EnVI algorithm falls under the NMF category (see definition in Appendix C), given the exploitation of the dependencies between latent states and the transition function, as evident in the variational distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ in Eq. (24). This indicates that EnVI inherits the favorable characteristics of NMF algorithms [27], which have the potential to enhance learning accuracy and address the issue of underestimating state inference uncertainty, commonly encountered in MF algorithms [33].
@@
Last but not least, our method offers an effective means of mitigating the risk of overfitting. This is mainly due to the fact that our method leverages the Bayesian nonparametric model and the variational inference framework to derive the ELBO, as presented in Eq. (26), from which the additional regularization terms for the initial state and state transition function can effectively mitigate the overfitting issue. As a comparison, in the state-of-the-art EnKF-based dynamical system learning method, autodifferentiable EnKF (AD-EnKF) [47, 51], the transition function in the SSM is modeled using a deterministic parametric model, specifically a neural network, and the optimization objective function is the logarithm of marginal likelihood of the model, i.e.
@@
Maximizing this objective function solely w.r.t. the model parameters can easily lead to overfitting. We defer further discussions of this issue to Section VI.
V OEnVI: Online Implementation of EnVI
In this section, we further explore online setting where data is processed sequentially, one sample at a time. It is within this context that the simultaneous inference of states and nonlinear dynamics in GPSSMs presents significant challenges [54, 55, 56]. The good news is that our EnVI approach readily lends itself to online learning scenarios. At each time step $t$, we can naturally maximize the corresponding objective function, denoted as $L_{\ell_{t}}$, given by
@@ V OEnVI: Online Implementation of EnVI
in terms of both the model parameters and variational parameters. Detailed steps for implementing the online EnVI, termed OEnVI, is summarized in Algorithm 2.
@@
Remark 3.
Proposition 2.
The log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, essentially is the difference between the data reconstruction error, represented by $\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]$, and the KL divergence between the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$ and the prediction distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$. That is,
$\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ (30)
$\!=\!\mathbb{E}_{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})}[\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})]\!-\!\operatorname{KL}\left[q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right].$
Thus, an alternative objective function for OEnVI can be expressed as
$\quad-\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\operatorname{KL}\left[q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\|p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]\right]$
Proof.
The proof can be found in Appendix B
This insight sheds light on the interplay between data reconstruction and the alignment of filtering and prediction distributions in the EnVI and OEnVI algorithms.
Up to this point, it is worth noting that in contrast to the existing inference network-based variational algorithms [26, 27, 28, 29, 35, 39, 56, 55], OEnVI is a simple and straightforward extension of EnVI, which benefits from eliminating the dependence on the additional parametric variational distributions. Previous works have typically employed inference networks that take the entire sequence of observations ${\mathbf{y}}_{1:T}$ as input, necessitating a significant amount of data for offline training. While it is conceivable to limit the input length of the inference network to a shorter sequence $l$, such as ${\mathbf{y}}_{t-l:t}$, this approach still leads to prolonged training times and higher computational requirements for optimizing the inference network parameters [56]. Moreover, storing historical inputs ${\mathbf{y}}_{t-l:t}$ adds to the storage overhead, making it problematic in situations where historical data duplication and storage are not permissible. In sharp contrast, OEnVI successfully overcomes the aforementioned challenges related to the optimization of inference networks. As a result, it facilitates more efficient learning and inference processes, contributing to potential improved overall performance. Furthermore, OEnVI offers a more principled objective, see Eq. (31), by simultaneously minimizing the KL divergence, accounting for data reconstruction error balance, and applying regularization to the transition function to mitigate model overfitting.
@@ VI-A Learning and Inference in Linear Gaussian SSMs
VI-B System Dynamics Learning
@@
Figure 4: Kink function learning performance against the training iterations. EnVI exhibits much faster convergence compared to vGPSSM and VCDT.
@@ VI-C Time Series Data Forecasting
This subsection demonstrates the series prediction performance of the proposed EnVI algorithm on five public real-world system identification datasets [^4] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html, which consist of one-dimensional time series of varying lengths between $296$ to $1024$ data points. In addition to the comparison methods discussed in Section VI-B, EnVI is also compared with several other competitors, including two NMF class methods, PRSSM [27], ODGPSSM [36], and two inference network-based methods, DKF [39] and CO-GPSSM [35], as depicted in Table II. For each method, the first half of the sequence in every dataset is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence, and the latent state dimension, $d_{x}$, is consistently set to $4$ for all datasets. Table II reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
@@
VI-D Online Learning and Inference Using OEnVI
Figs. 5a5c depict the true states (in blue) and the latent states (in red) inferred by the three methods. The results clearly indicate that EnVI and SVMC swiftly captured the real state and maintained accuracy, while VJF faced challenges, primarily due to its difficulty in optimizing the parameters in the inference network. The detailed comparison with SVMC, illustrated in Figs. 5d5f, highlights the superior accuracy of OEnVI in both inference and prediction of latent states. These empirical findings emphasize the efficacy of OEnVI, particularly in its use of EnKF to approximate the variational distribution, demonstrating its advancement over VJF using inference networks, and SVMC using PF methods.
VII Conclusion
In this paper, we have introduced EnVI, a novel NMF algorithm tailored for GPSSMs, which integrates EnKF into a variational inference framework. Additionally, we have presented an extended online version, OEnVI. Both algorithms eliminate the necessity for heavy parameterization like inference networks and shape the variational distribution over latent states through the model-based EnKF. Leveraging the inherent differentiable nature along with the modern automatic differentiation tools, the proposed EnVI and OEnVI can enhance efficiency and algorithmic robustness while improving learning and inference performance compared to existing methods. Detailed analysis and fresh insights for the proposed algorithms are provided to enhance their interpretability. Empirical experiments conducted on diverse real and synthetic datasets, evaluating filtering, prediction, and dynamics learning performance, unequivocally support the effectiveness of the proposed methods.
Appendix A Model Evidence Lower Bound Evaluations
With the model joint distribution, Eq. (16), and the variational distribution, Eq. (19), we can write down the ELBO according to the general definition given in Eq. (7), i.e.,
where the KL divergence terms can be analytically computed in closed form [10] due to the Gaussian nature of the prior and variational distributions. The evaluation of the first and fourth terms is typically intractable. We examine the difference between term 1 and term 4 in Eq. (36) and can have the following lemma.
Lemma 1.
Under the approximations that $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, $\forall t$, computing the difference between term 1 and term 4 in the ELBO (Eq. (21)) yields an approximation to an expected log-likelihood, i.e.,
$\operatorname{term~{}1}-\operatorname{term~{}4}\approx\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{1:T}|{\vec{\mathbf{u}}})\right]$ (37)
@@ Appendix A Model Evidence Lower Bound Evaluations
According to the ELBO given in Eq. (36), we have:
@@
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right],$ (38e)
@@
According to Lemma 1, and the ELBO given in Eq. (36), we immediately get the following approximated ELBO:
$\!-\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})],$ (40)
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (39) can be analytically evaluated, see Eq. (27), due to the Gaussian prediction distribution, see Eq. (23), and the linear emission model.
@@ Appendix B Proof of Proposition 2
which completes the proof. Here Eq. (41c) is obtained straightforwardly by applying Bayes’ theorem. This result sheds light on the interplay between data reconstruction and the alignment of filtering and prediction distributions in the EnVI and OEnVI algorithms. ∎
@@ D-A Evidence Lower Bound (ELBO)
In Bayesian statistics, the model marginal likelihood $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is a fundamental quantity for model selection and comparison [38]. By maximizing the logarithm of $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ w.r.t. the model parameters ${{\bm{\theta}}}$, the goodness of data fitting and the model complexity are automatically balanced, in accordance with Occam’s razor principle [37]. However, $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is obtained by integrating out all the latent variables $\{{\vec{\mathbf{x}}},{\vec{\mathbf{f}}}\}$ in the joint distribution, see Eq. (15), which is analytically intractable. Thus, the posterior distribution of the latent variables, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})=\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{y}}}|{{\bm{\theta}}})},$ cannot be expressed in a closed-form expression, either. This intractability issue has been addressed in variational Bayesian methods by adopting a variational distribution [10], $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, to approximate the intractable $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$. With the newly introduced variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, we have